ABCD is a rectangle. The area of the rectangle is 40 square feet.
The length is 3 more than the width.
A) Find the measure of each length and each width
B) Find the length of the diagonal, rounded to the nearest tenth

Answer by mananth(16946)

Area of rectangle = l*w
Width be x
Length = x+3
x(x+3) =40
x^2+3x=40
x^2+3x-40=0
x^2+8x-5x-40=0
x(x+8)-5(x+8)=0
(x+8)(x-5)=0
Therefore x= -8 OR x=5
Ignore negative
x=5 is the width
width = 5 feet

Length = x+3----> 8feet